OpenStudy (anonymous):

The Sugar Sweet Company will choose from two companies to transport its sugar to market. The first company charges $7500 to rent trucks plus an additional fee of $125.25 for each ton of sugar. The second company charges $5995 to rent trucks plus an additional fee of $200.50 for each ton of sugar.

OpenStudy (anonymous):

for what amount of sugar do the two companies charge the same?

OpenStudy (anonymous):

in ton

OpenStudy (anonymous):

Ok you have two equations here, one from the first company and one from the second company. Do you know what it is?

OpenStudy (anonymous):

no i dont know the equations

OpenStudy (anonymous):

Ok no problem I'll show you how I got the first equation, So the first company has a initial charge of $7500 plus $125.25 per ton of sugar. If we put it into an equation where x is number of tons in sugar, you'll get y=7500+125.25x Now, do you think you can do the second one?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

just replace the numbers for the second one

OpenStudy (anonymous):

y=5995+200.50x

OpenStudy (anonymous):

Ok great, set the two equations equal to each other to find x. To find when the amount of sugar is charged equally by the two companies. 7500+125.25x=5995+200.50x Or if you have a graphing calculator just graph it and find the intersecting point.

OpenStudy (anonymous):

so solve both together?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

Yep that's correct just make sure to include tons for your unit.

OpenStudy (anonymous):

If you are unsure if its correct, just plug it back in two both equations and you should get the same answer.

OpenStudy (anonymous):

there is another questin for it

OpenStudy (anonymous):

what is the cost when the two companies charge the same?

OpenStudy (anonymous):

Then plug 20 back into x for the two equations. You'll get the charge price.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

7500+125.25x=5995+200.50x instead of x i put 20?

OpenStudy (anonymous):

No, plug 20 in for y=7500+125.25x and y=5995+200.5x

OpenStudy (anonymous):

plug 20 in for x and you'll get total price for y

OpenStudy (anonymous):

i solve each?

OpenStudy (anonymous):

You can if you want, or just choose one because they'll both equal the same charge price.

OpenStudy (anonymous):

\[y=7500+125.25(20)=7500+2505=10005\]\[y=5995+200.5(20)=5995+4010=10005\]

OpenStudy (anonymous):

yes what i got when evaluating

OpenStudy (anonymous):

so that is the cost 1005?

OpenStudy (anonymous):

Ok so 10005 is the amount the two companies charge, $10,005

OpenStudy (anonymous):

Thank you very much

OpenStudy (anonymous):

No problem, hope it helps and good luck :)